You intend to conduct an ANOVA with 7 groups in which each group will have the same number of subjects: n=6n=6. (This is referred to as a "balanced" single-factor ANOVA.)

What are the degrees of freedom for the numerator?
d.f.(treatment) =

What are the degrees of freedom for the denominator?
d.f.(error) =
Answer #1

Solution:

Given k=7, n=6;

Total df= nk-1= 7*6-1=42-1=41

Treatment df= k-1=7-1=6

Error df= Total df- Treatment df =41-6=35

Thus, d.f (treatment) = 6

d.f (error) =35

Source of Variation Degrees of freedom
Treatment k-1=7-1=6
Error Total df- Treat df= 41-6 =35
Total nk-1= 7*6-1= 42-1=41